# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

### v3

@@ 4 Experimental Setup
  • FinCoT: A zero-shot prompting method that integrates expert domain templates. Each prompt includes a Mermaid diagram as a “Hint” to guide structured financial reasoning via relevant domain insights.
@@ 5.1 Financial Reasoning Performance
On Qwen2.5-7B, FinCoT (All Blueprints) yields a +7.95 pp improvement over SP (95% CI [6.30, 9.59], $p<0.001$), while UST-CoT and ST-CoT perform worse than the baseline. When FinCoT is applied using a single-domain blueprint (e.g., Financial Reporting), the gains increase substantially to +15.21 pp. Similarly, on Qwen3-8B-Base, FinCoT delivers the strongest overall boost (+17.35 pp, 95% CI [15.02, 19.77], $p<0.001$). These findings further underscore the importance of structured domain knowledge, particularly for models lacking instruction or domain alignment.
@@
This section examines how pretrained and finance-specific models respond to structured prompting with FinCoT. Each domain-specific blueprint is applied across all CFA domains to evaluate its transferability, and accuracy differences relative to SP are measured. Figure 4(a) and 4(b) visualize results for Qwen3-8B-Base (pretrained) and Fin-o1-8B (finance-specific), while Table 1 provides a comprehensive summary of overall model-level accuracy. On Qwen3-8B-Base, FinCoT consistently improves performance. Prompts from quantitative domains such as Derivatives, Portfolio Management, and Corporate Issuers yield average gains exceeding +13 pp. The blueprint structure provides inductive guidance that enhances decomposition, formula selection, and financial term alignment. In several domains, Qwen3-8B-Base with FinCoT matches or surpasses the SP baseline of Fin-o1-8B, despite the absence of any task-specific training.
@@ 5.2 Efficiency Analysis
Effective deployment of foundation models in financial settings requires not only accuracy but also efficiency in output length and computational cost. This section investigates the verbosity and accuracy trade-offs across prompting strategies. While Fin-o1-8B, DianJin-R1-7B, and Fin-R1 (7B) are fine-tuned models optimized for financial reasoning, FinCoT offers a prompting-based alternative that achieves competitive performance without additional training. Comparative accuracy is illustrated in Figure 5, and a detailed breakdown of output lengths across models and domains is provided in Appendix D (Table 2) to support efficiency analysis.
@@
In practical deployments, prompting strategies that achieve high accuracy with minimal output length are especially desirable. This section compares the token efficiency of FinCoT with standard prompting strategies, focusing on general-purpose and financial-specific models.
@@
Three behavioral zones emerge from these results. First, models like Gemma-3-12B-IT and Qwen3-8B-Base show high mean improvement and strong cross-domain transfer, reflecting effective generalization. Second, models such as Qwen2.5-7B-Instruct and Qwen3-8B display noticeable cross-domain transfer with lower mean improvement, suggesting limited benefit. Third, models like Fin-o1-8B and Qwen3-8B (Thinker) exhibit low cross-domain adaptability and minimal performance lift, with signs that overly detailed prompting may interfere with internal reasoning.
These findings highlight the importance of aligning the prompting strategy with a model’s pretraining or fine-tuning characteristics to optimize both performance and efficiency. With limited supervision, FinCoT offers a transparent, cost-effective alternative for enhancing financial reasoning.
@@ 6 Conclusion
(i) Blueprint creation currently requires expert review ($\sim$2 hours/domain); (ii) evaluations are limited to multiple-choice settings; (iii) interpretability is measured using rubric scores, not human studies.
Future work includes automating blueprint generation and integrating retrieval-augmented diagrams to expand FinCoT’s applicability.
Our findings suggest that expert-guided prompting acts as a lightweight form of knowledge distillation, embedding structured reasoning in LLM outputs without altering model weights. This method may generalize to high-stakes fields like law, medicine, and engineering, and underscores the value of prompt-level supervision for interpretable, efficient domain reasoning.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Credit Risk Assessment and Rating Measures – Evaluate creditworthiness through agency ratings (e.g., S&P, Moody’s), credit spreads, and financial ratios. Analyze probability of default and credit cycle indicators. This step is vital for bondholders and fixed income portfolio managers (Fabozzi, 2012).
@@ Appendix C Domain Distribution
Appendix D Average Output Tokens

### v4

@@ 4 Experimental Setup
  • FinCoT: A zero-shot prompting method that integrates expert domain templates (excluding the Ethics domain). Each prompt includes a Mermaid diagram as a “Hint” to guide structured financial reasoning via relevant domain insights.
@@ 5.1 Financial Reasoning Performance
On Qwen2.5-7B, FinCoT (All Blueprints) yields a +7.95 pp improvement over SP (95% CI [6.30, 9.59], $p<0.001$), while UST-CoT and ST-CoT also exceed the baseline by +13.76 pp and +16.28 pp. When FinCoT is applied using a single-domain blueprint (e.g., Financial Reporting), the gains increase substantially to +15.21 pp. Similarly, on Qwen3-8B-Base, FinCoT delivers the strongest overall boost (+17.35 pp, 95% CI [15.02, 19.77], $p<0.001$). These findings further underscore the importance of structured domain knowledge, particularly for models lacking instruction or domain alignment.
@@
This section examines how pretrained and finance-specific models respond to structured prompting with FinCoT. Each domain-specific blueprint is applied across all CFA domains to evaluate its transferability, and accuracy differences relative to SP are measured. Figure 4(a) and 4(b) visualize results for Qwen3-8B-Base (pretrained) and Fin-o1-8B (finance-specific), while Table 1 provides a comprehensive summary of overall model-level accuracy. On Qwen3-8B-Base, FinCoT generally improves performance, though gains are not universal. Prompts from quantitative domains such as Derivatives, Portfolio Management, and Corporate Issuers yield average gains exceeding +13 pp. The blueprint structure provides inductive guidance that enhances decomposition, formula selection, and financial term alignment. In several domains, Qwen3-8B-Base with FinCoT matches or surpasses the SP baseline of Fin-o1-8B, despite the absence of any task-specific training.
@@ 5.2 Efficiency Analysis
Effective deployment of foundation models in financial settings requires balancing verbosity and accuracy while emphasizing efficiency. FinCoT offers a prompt-based alternative to fine-tuned models (Fin-o1, DianJin-R1, Fin-R1), demonstrating similar accuracy as seen in Fig. 5, with token lengths detailed in Appendix D.2 (Tab. 3). Our analysis concentrates on output tokens since prompt encoding occurs once with parallel self-attention $O(n_{\text{in}}^{2})$ (rapid) (Vaswani et al., 2023), while decoding involves $O(n_{\text{out}})$ sequential steps with per-step KV-cache updates (high memory usage) (Gu et al., 2018; Shazeer, 2019). Comprehensive input and output token data are provided in Appendix D.1(Tab. 2), and recent decoding accelerations such as speculative sampling (Chen et al., 2023) and FlashAttention (Dao et al., 2022) further underscore output as the primary latency driver.
@@
In deployments, prompting strategies that achieve high accuracy with minimal output length are desirable.
@@
Three behavioral zones emerge from these results. First, models like Gemma-3-12B-IT and Qwen3-8B-Base show high mean improvement and strong cross-domain transfer, reflecting effective generalization. Second, models such as Qwen2.5-7B-Instruct and Qwen3-8B display noticeable cross-domain transfer with lower mean improvement, suggesting limited benefit. Third, models like Fin-o1-8B and Qwen3-8B (Thinker) exhibit low cross-domain adaptability and minimal performance lift, indicating that overly detailed prompting may conflict with internal reasoning.
These findings underscore the importance of aligning the prompting strategy with a model’s pretraining or fine-tuning to optimize performance and efficiency. With limited supervision, FinCoT provides a transparent, cost-effective alternative for enhancing financial reasoning. For a price–sensitivity analysis of token cost, see Appendix D.2.1 and Fig. 7.
@@ 6 Conclusion
Our evaluation highlights several limitations. (i) Efficiency gains come mainly from reduced output tokens, but larger inputs from structured templates add cost; overall, FinCoT is still competitive, especially in long-form reasoning. (ii) Domain routing via pre-classification risks template mismatch despite safeguards; adaptive selection methods are needed. (iii) Improvements are uneven, with domains like Alternative Investments and Ethics limited by small samples ($\sim$10–30); larger, balanced benchmarks are required. (iv) Blueprint creation requires expert effort ($\sim$2 hours/domain), and current evaluations are multiple-choice with rubric-based interpretability.
Overall, FinCoT offers structured, auditable reasoning rather than replacing fine-tuned models. Its blueprint methodology and plug-and-play usability demonstrate prompt-level supervision as lightweight knowledge distillation with potential for law, medicine, and engineering.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Credit Risk Assessment and Rating Measures – Evaluate creditworthiness through agency ratings (e.g., S&P, Moody’s), credit spreads, and financial ratios. Analyze the probability of default and credit cycle indicators. This step is vital for bondholders and fixed income portfolio managers (Fabozzi, 2012).
@@ Appendix C Domain Distribution
Appendix D Average Input and Output Tokens
D.1 Average Input Tokens
Prompt Average Input Tokens (k)
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Gemma-3-12B
IT
Qwen3-8B
Fin-R1
DianJin-R1
7B
Fin-o1-8B
SP 0.07* 0.07* 0.07* 0.07* 0.07* 0.07* 0.07* 0.07* 0.07*
UST-CoT 0.09 0.09 0.09 0.09 0.09 0.09 0.09 0.09 0.09
ST-CoT 0.18 0.18 0.18 0.18 0.19 0.18 0.18 0.18 0.18
FinCoT (All Blueprints) 1.75 1.75 1.75 1.75 1.78 1.75 1.75 1.75 1.75
Domain-wise performance of FinCoT
FinCoT (Economics) 0.55 0.55 0.55 0.55 0.56 0.55 0.55 0.55 0.55
FinCoT (FixedIncome) 0.34 0.34 0.34 0.34 0.36 0.34 0.34 0.34 0.34
FinCoT (Quant.Meth.) 0.33 0.33 0.33 0.33 0.34 0.33 0.33 0.33 0.33
FinCoT (EquityInvest.) 0.34 0.34 0.34 0.34 0.36 0.34 0.34 0.34 0.34
FinCoT (Port.Manage.) 0.33 0.33 0.33 0.33 0.35 0.33 0.33 0.33 0.33
FinCoT (Derivatives) 0.39 0.39 0.39 0.39 0.44 0.39 0.39 0.39 0.39
FinCoT (Fin. Reporting) 0.37 0.37 0.37 0.37 0.38 0.37 0.37 0.37 0.37
FinCoT (Alter.Invest.) 0.32 0.32 0.32 0.32 0.34 0.32 0.32 0.32 0.32
FinCoT (Corp.Issuers) 0.39 0.39 0.39 0.39 0.41 0.39 0.39 0.39 0.39
Table 2: Comparison of prompting techniques: average input token length (k) across models. Bold values highlight the prompt variant that uses the least tokens for each model.
D.2 Average Output Tokens
@@
D.2.1 Efficiency of Input and Output Cost in Simulation
This appendix reports a cost–efficiency analysis under realistic output–input price ratios. Let $I$ and $O$ denote the average input and output tokens for a (prompt, model) pair. For a price ratio $r$,
$\mathrm{Cost}(r)=I+r\,O,\qquad\mathrm{Efficiency}(r)=\frac{\mathrm{Cost}_{\mathrm{baseline}}(r)}{\mathrm{Cost}_{\mathrm{prompt}}(r)}=\frac{I_{\mathrm{base}}+r\,O_{\mathrm{base}}}{I_{\mathrm{prompt}}+r\,O_{\mathrm{prompt}}}.$
Units and normalization. We measure cost in “input-token dollars”: the effective input price is $1$, and $r=\text{price}_{\text{out}}/\text{price}_{\text{in,eff}}$ carries the output premium. This rescaling makes $\mathrm{Efficiency}$ dimensionless and invariant to any common price factor.
Break-even and sensitivity. For a candidate prompt $p$ vs. baseline $b$, the break-even ratio solving $\mathrm{Cost}_{p}(r)=\mathrm{Cost}_{b}(r)$ is
$r^{\star}=\frac{I_{p}-I_{b}}{\,O_{b}-O_{p}\,}\qquad(O_{p}\neq O_{b}).$
If $O_{p}<O_{b}$, then $\dfrac{d}{dr}\mathrm{Efficiency}(r)=\dfrac{O_{b}I_{p}-O_{p}I_{b}}{(I_{p}+rO_{p})^{2}}>0$: the candidate improves as $r$ increases; if $O_{p}>O_{b}$, the trend reverses. When $O_{p}=O_{b}$, ranking depends only on inputs ($I_{b}$ vs. $I_{p}$) and is independent of $r$.
Caching and effective input price. With prompt caching,
$p_{\mathrm{in,eff}}(K)=p_{\mathrm{read}}+\frac{p_{\mathrm{write}}}{K},\qquad r(K)=\frac{\text{price}_{\mathrm{out}}}{p_{\mathrm{in,eff}}(K)},$
where $K$ is the number of reuses. Hence $r(K)$ increases monotonically in $K$ and approaches $\text{price}_{\mathrm{out}}/p_{\mathrm{read}}$ as $K\!\to\!\infty$.
[…]
Worked example (illustrative). Baseline $(I_{b},O_{b})=(100,300)$; candidate $(I_{p},O_{p})=(250,150)$. At $r=8$ (GPT-5 no cache): $\mathrm{Cost}_{b}=100+8\!\cdot\!300=2500$, $\mathrm{Cost}_{p}=250+8\!\cdot\!150=1450$, so $\mathrm{Efficiency}\approx 1.72$. Break-even $r^{\star}=(250-100)/(300-150)=1$; the candidate dominates for $r>1$.
Note (scope). All models evaluated in this appendix are open-source. The curves simulate dollar costs by pairing the measured $(I,O)$ token counts from these models with provider API prices (GPT-5 and Claude Opus 4.1)—prices are used for simulation only; no paid API runs were executed for these experiments.
Figure 7: Cost–efficiency vs. price ratio $r$ for UST-CoT, ST-CoT, and FinCoT-All across models. Efficiency is $\mathrm{Cost}_{\mathrm{baseline}}/\mathrm{Cost}_{\mathrm{prompt}}$ with $\mathrm{Cost}=I+rO$; values >1>1 indicate lower cost than the baseline. $r$ values use provider prices for GPT-5 and Claude Opus 4.1 as described above. Notation: MTok = million tokens; USD per MTok.